OpenStudy (anonymous):

Hi ( :can anyone help me solve this? Use interval notation to state the range of the quadratic function. f(x) = -2(x + 1)^2 - 3

OpenStudy (anonymous):

lets take it step by step

OpenStudy (anonymous):

the smallest \((x+1)^2\) can be is 0, because if you square any non zero number it is positive

OpenStudy (anonymous):

therefore the LARGEST \(-2(x+1)^2\) can be is zero, otherwise it is negative (less than zero)

OpenStudy (anonymous):

and finally that means the largest \[-2(x+1)^2-3\] can be is \(-3\)

OpenStudy (anonymous):

that makes the range all numbers less than or equal to \(-3\) which, in interval notation, is \[(-\infty, -3]\)

OpenStudy (anonymous):

\[(-\infty, -3]\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

would please help me with another question?

OpenStudy (anonymous):

On which interval is f(x) = (x - 6)2 + 3 decreasing?